This question is from textbook college algebra
: It is approximately 300 miles from chicago,illinois to st. louis,missouri. allowing for various traffic conditions, a driver can average approximately 60 miles per hour.
d) how far will i need to travel to rerah St.Louis after i have traveled 3 hours.
Thank you for the help in advance.
textbook info.
ISBN 0-321-27908-5
TITLE
COLLEGE ALGEBRA BY GARY ROCKSWOLD
3RD EDITION
PROBLEM 3

Answer by Nate(3500)

300 - 60x = d |where x is the hours driven and d is the distance between destinations
300 - 60(3) = d
300 - 180 = d
120 = d